Solve 45x^2-60x+20=0: Solve 45x^2-60x+20=0
Answer by jim_thompson5910(9403)

45x^2-60x+20 Start with the given expression


5(9x^2-12x+4) Factor out the GCF 5


Now let's focus on the inner expression 9x^2-12x+4


------------------------------------------------------------

Looking at the expression 9x^2-12x+4, we can see that the first coefficient is 9, the second coefficient is -12, and the last term is 4.


Now multiply the first coefficient 9 by the last term 4 to get (9)(4)=36.


Now the question is: what two whole numbers multiply to 36 (the previous product) and add to the second coefficient -12?


To find these two numbers, we need to list all of the factors of 36 (the previous product).


Factors of 36:
1,2,3,4,6,9,12,18,36
-1,-2,-3,-4,-6,-9,-12,-18,-36


Note: list the negative of each factor. This will allow us to find all possible combinations.


These factors pair up and multiply to 36.
1*36
2*18
3*12
4*9
6*6
(-1)*(-36)
(-2)*(-18)
(-3)*(-12)
(-4)*(-9)
(-6)*(-6)

Now let's add up each pair of factors to see if one pair adds to the middle coefficient -12:


First NumberSecond NumberSum
1361+36=37
2182+18=20
3123+12=15
494+9=13
666+6=12
-1-36-1+(-36)=-37
-2-18-2+(-18)=-20
-3-12-3+(-12)=-15
-4-9-4+(-9)=-13
-6-6-6+(-6)=-12



From the table, we can see that the two numbers -6 and -6 add to -12 (the middle coefficient).


So the two numbers -6 and -6 both multiply to 36 and add to -12


Now replace the middle term -12x with -6x-6x. Remember, -6 and -6 add to -12. So this shows us that -6x-6x=-12x.


9x^2+highlight(-6x-6x)+4 Replace the second term -12x with -6x-6x.


(9x^2-6x)+(-6x+4) Group the terms into two pairs.


3x(3x-2)+(-6x+4) Factor out the GCF 3x from the first group.


3x(3x-2)-2(3x-2) Factor out 2 from the second group. The goal of this step is to make the terms in the second parenthesis equal to the terms in the first parenthesis.


(3x-2)(3x-2) Combine like terms. Or factor out the common term 3x-2


(3x-2)^2 Condense the common factors.


So the expression goes from 5(9x^2-12x+4) and factors further to 5(3x-2)^2




So 45x^2-60x+20 factors to 5(3x-2)^2.



5(3x-2)^2=0 Set the factored expression equal to zero


(3x-2)^2=0 Divide both sides by 5.


3x-2=0 Take the square root of both sides


3x=0+2 Add 2 to both sides.


3x=2 Combine like terms on the right side.


x=(2)/(3) Divide both sides by 3 to isolate x.


----------------------------------------------------------------------

Answer:

So the answer is x=2/3


Which approximates to x=0.667